# Problem Solved? Information Extraction Design Space for Layout-Rich Documents using LLMs

### v3

@@ 1 Introduction
Information extraction (IE) extracts structured data from unstructured documents, including layout-rich documents (LRDs) as reports and presentations that mix visual and textual elements Park et al. (2019); Wang et al. (2023b); Zmigrod et al. (2024b). LRDs challenge traditional natural language processing (NLP) techniques, designed for plain texts Cui et al. (2021); Tang et al. (2023).
@@
Recent layout-aware models at the intersection of NLP and Computer Vision (CV) address this gap by including visual and structural features to improve IE from LRDs (e.g., LayoutLMv1-v3 Xu et al. (2020b, a); Huang et al. (2022), ERNIE-Layout Peng et al. (2022), GraphDoc Zhang et al. (2022), DocFormer Appalaraju et al. (2021)). While early layout-aware models required substantial dataset-specific fine-tuning, recent generative document understanding models Powalski et al. (2021) have demonstrated strong zero- and few-shot capabilities on unseen datasets and tasks.
@@
Our results reveal multiple insights. First, current general LLMs can compete with traditional fine-tuned models such as LayoutLMv3 and ERNIE-Layout, without expensive labeling. Second, rather than fine-tuning on data, LLMs require adjustment of the IE pipeline to achieve competitive performance–the performance gap between a general-practice baseline and the configuration tuned using our lightweight OFAT method on two datasets is 13.3 F1 points for VRDU Wang et al. (2023c), and 37.5 for FUNSD Jaume et al. (2019), only 1.8 points and 0.8 points lower than the best possible Brute-Force configuration respectively. Third, while purely text-based LLMs achieve competitive performance with our method, multimodal LLMs, those that directly integrate textual and visual features, still outperform them. However, this advantage comes with increased token usage, higher API costs, and diminished transparency and control over individual steps. In summary, our contributions are as follows.
@@ 2.1 Using LLMs for IE
Model Engagement. Once preprocessed, the document is fed to an LLM for IE. Ensuring alignment between the extracted text and layout information is crucial for accurate representation Xu et al. (2020a); Appalaraju et al. (2021). Prompt-driven extraction leverages general-purpose models, by using tailored prompts to guide the extraction process Brown et al. (2020); Radford et al. (2021); Zhou et al. (2022). As such, models must be explicitly instructed to extract information, typically with an accompanying schema. This step can involve more advanced IT and ICL techniques (few-shot, CoT). The influence of prompting techniques in interaction with various stages of the IE pipeline to enhance performance and robustness remains a research gap that requires further investigation.
Output Refinement. After inference, the extracted information undergoes post-processing to ensure accuracy and conformity with Schema $S$. This step involves refining and validating the outputs generated by the LLM through tasks such as mapping extracted entities $E$ to their original document positions, merging overlapping or fragmented predictions, and resolving ambiguities in the results Xu et al. (2020b). Post-processing for entity extraction has been explored in prior studies Wang et al. (2022b); Tamayo et al. (2022), and rule-based entity alignment has demonstrated notable improvements in accuracy Luo et al. (2024b). However, to our knowledge, there has been no analysis of post-processing techniques specifically tailored to LRDs in conjunction with LLMs.
@@ 3.3 Output Refinement
We refine the raw outputs from the LLM to ensure alignment with the target schema, addressing challenges related to prediction variability, schema definition differences, and data formatting inconsistencies (e.g., varying date formats). We implement three techniques inspired by related work in data integration Dong and Srivastava (2013): Decoding, Schema Mapping, and Data Cleaning. This results in three sets of predictions: initial predictions, mapped predictions, and cleaned predictions.
@@ 4.1 Experimental Design
Dataset and LLMs. We use the Visually Rich Document Understanding (VRDU) dataset Wang et al. (2023c), which includes two benchmarks with high-quality OCR for assessing data efficiency, due to its inclusion of three visually diverse template types – Amendment, Short-Form, and Dissemination – representing varied and complex layouts, as well as generalization tasks: Single Template Learning (STL), Unseen Template Learning (UTL), and Mixed Template Learning (MTL). For further details on how this dataset is tailored for our experiments and diverse models, please refer to A.3.
@@ 5 Related Work
IE using LLMs. Transformer-based models Devlin et al. (2019); Brown et al. (2020); Liu et al. (2019b) have advanced NLP with self-attention and large-scale pretraining but struggle with layout-rich documents (LRDs). To address this, layout-aware and multimodal models have emerged Xu et al. (2020b, 2021, 2022), including generative approaches Powalski et al. (2021). Additionally, some models enhance document understanding through multimodal learning OpenAI (2023); Anil et al. (2023). Structural-aware approaches Fujitake (2024); Lee et al. (2022) further improve extraction accuracy, while end-to-end models Kim et al. (2022) bypass OCR for direct document image processing. Furthermore, some approaches adopt a chat-based paradigm for flexible IE Xu et al. (2024a), while others Jiang et al. (2025) enhance document understanding through layout-aware pretraining, advancing LLMs for practical applications.
Strategies for IE from LRDs. Graph-based models Liu et al. (2019a); Nourbakhsh et al. (2024) enhance relation extraction by capturing textual-visual relationships. A critical factor in layout-rich document (LRD) processing is maintaining the reading order, where Token Path Prediction (TPP)Zhang et al. (2023) resolves OCR layout ambiguities, and global taggingShaojie et al. (2023) mitigates text ordering issues, thereby improving extraction accuracy. For structured data, specific approaches target various document types: table extraction Jain et al. (2020), form processing Zhong et al. (2020), and unified OCR, preprocessing, and postprocessing for document IE Perot et al. (2024). Additionally, a simple yet effective multimodal and multilingual method for parsing semi-structured forms is presented in Cheng et al. (2025). Despite the variety of models proposed, comparative studies evaluating text-only, multimodal, and layout-aware LLMs remain limited Van Landeghem et al. (2023); Stanisławek et al. (2021).
Preprocessing, Chunking, Prompting, Postprocessing, and Evaluation Techniques. Chain-of-Thought (CoT) prompting Wei et al. (2022) enhances reasoning in complex LRD extraction tasks, while diverse prompt-response datasets Zmigrod et al. (2024a) improve LLM robustness. Instruction-finetuned LLMs have also demonstrated effectiveness in domain-specific applications, such as clinical and biomedical tasks, where zero-shot and few-shot enable adaptive extraction without extensive fine-tuning Labrak et al. (2024). Postprocessing techniques, including text normalization, schema alignment Fürst et al. (2023), entity resolution Hwang et al. (2021), and majority voting Wang et al. (2022a), refine extracted data by correcting OCR and extraction errors.
@@ Acknowledgments
Our experimental evaluation focused on the Visually Rich Document Understanding (VRDU) dataset Wang et al. (2023c) and the FUNSD dataset Jaume et al. (2019). VRDU was chosen for its comprehensive OCR outputs, well-defined schema, and diverse template types (Amendment, Short-Form, and Dissemination) that capture complex layouts. It also supports generalization tasks–Single Template Learning (STL), Unseen Template Learning (UTL), and Mixed Template Learning (MTL)–making it suitable for visually rich document understanding. In contrast, FUNSD was selected due to its widespread use in document understanding tasks, particularly for extracting structured information from simpler, form-based layouts. While FUNSD does not share the layout diversity and complexity of VRDU, it serves as a useful baseline for form-centric document analysis. This provides a different perspective by offering insights into scenarios where key-value pairs are organized within relatively simple templates.

### v4

@@ 1 Introduction
Information extraction (IE) extracts structured data from unstructured documents, including layout-rich documents (LRDs) as reports and presentations that mix visual and textual elements Park et al. (2019); Wang et al. (2023b, 2025). LRDs challenge traditional natural language processing (NLP) techniques, designed for plain texts Cui et al. (2021); Tang et al. (2023).
@@
Recent layout-aware models at the intersection of NLP and Computer Vision (CV) address this gap by including visual and structural features to improve IE from LRDs (e.g., LayoutLMv1-v3 Xu et al. (2020, 2021); Huang et al. (2022), ERNIE-Layout Peng et al. (2022), GraphDoc Zhang et al. (2022), DocFormer Appalaraju et al. (2021)). While early layout-aware models required substantial dataset-specific fine-tuning, recent generative document understanding models Powalski et al. (2021) have demonstrated strong zero- and few-shot capabilities on unseen datasets and tasks.
@@
Our results reveal multiple insights. First, current general LLMs can compete with traditional fine-tuned models such as LayoutLMv3 and ERNIE-Layout, without expensive labeling. Second, rather than fine-tuning on data, LLMs require adjustment of the IE pipeline to achieve competitive performance–the performance gap between a general-practice baseline and the configuration tuned using our lightweight OFAT method on two datasets is 13.3 F1 points for VRDU Wang et al. (2023b), and 37.5 for FUNSD Jaume et al. (2019), only 1.8 points and 0.8 points lower than the best possible Brute-Force configuration respectively. Third, while purely text-based LLMs achieve competitive performance with our method, multimodal LLMs, those that directly integrate textual and visual features, still outperform them. However, this advantage comes with increased token usage, higher API costs, and diminished transparency and control over individual steps. In summary, our contributions are as follows.
@@ 2.1 Using LLMs for IE
Model Engagement. Once preprocessed, the document is fed to an LLM for IE. Ensuring alignment between the extracted text and layout information is crucial for accurate representation Xu et al. (2021); Appalaraju et al. (2021). Prompt-driven extraction leverages general-purpose models, by using tailored prompts to guide the extraction process Brown et al. (2020); Radford et al. (2021); Zhou et al. (2022). As such, models must be explicitly instructed to extract information, typically with an accompanying schema. This step can involve more advanced IT and ICL techniques (few-shot, CoT). The influence of prompting techniques in interaction with various stages of the IE pipeline to enhance performance and robustness remains a research gap that requires further investigation.
Output Refinement. After inference, the extracted information undergoes post-processing to ensure accuracy and conformity with Schema $S$. This step involves refining and validating the outputs generated by the LLM through tasks such as mapping extracted entities $E$ to their original document positions, merging overlapping or fragmented predictions, and resolving ambiguities in the results Xu et al. (2020). Post-processing for entity extraction has been explored in prior studies Wang et al. (2022); Tamayo et al. (2022), and rule-based entity alignment has demonstrated notable improvements in accuracy Luo et al. (2024b). However, to our knowledge, there has been no analysis of post-processing techniques specifically tailored to LRDs in conjunction with LLMs.
@@ 3.3 Output Refinement
We refine the raw outputs from the LLM to ensure alignment with the target schema, addressing challenges related to prediction variability, schema definition differences, and data formatting inconsistencies (e.g., varying date formats). We implement three techniques inspired by related work in data integration Srivastava and Dong (2013): Decoding, Schema Mapping, and Data Cleaning. This results in three sets of predictions: initial predictions, mapped predictions, and cleaned predictions.
@@ 4.1 Experimental Design
Dataset and LLMs. We use the Visually Rich Document Understanding (VRDU) dataset Wang et al. (2023b), which includes two benchmarks with high-quality OCR for assessing data efficiency, due to its inclusion of three visually diverse template types – Amendment, Short-Form, and Dissemination – representing varied and complex layouts, as well as generalization tasks: Single Template Learning (STL), Unseen Template Learning (UTL), and Mixed Template Learning (MTL). For further details on how this dataset is tailored for our experiments and diverse models, please refer to A.3.
@@ 5 Related Work
IE using LLMs. Transformer-based models Devlin et al. (2019); Brown et al. (2020); Liu et al. (2019b) have advanced NLP with self-attention and large-scale pretraining but struggle with layout-rich documents (LRDs). To address this, layout-aware and multimodal models have emerged Xu et al. (2020, 2021); Huang et al. (2022), including generative approaches Powalski et al. (2021). Additionally, some models enhance document understanding through multimodal learning OpenAI (2023); Team et al. (2023). Structural-aware approaches Fujitake (2024); Lee et al. (2022) further improve extraction accuracy, while end-to-end models Kim et al. (2022) bypass OCR for direct document image processing. Furthermore, some approaches adopt a chat-based paradigm for flexible IE Xu et al. (2024a), while others Jiang et al. (2025) enhance document understanding through layout-aware pretraining, advancing LLMs for practical applications.
Strategies for IE from LRDs. Graph-based models Liu et al. (2019a); Nourbakhsh et al. (2024) enhance relation extraction by capturing textual-visual relationships. A critical factor in layout-rich document (LRD) processing is maintaining the reading order, where Token Path Prediction (TPP)Zhang et al. (2023) resolves OCR layout ambiguities, and global taggingShaojie et al. (2023) mitigates text ordering issues, thereby improving extraction accuracy. For structured data, specific approaches target various document types: table extraction Shigarov et al. (2018) and unified OCR, preprocessing, and postprocessing for document IE Perot et al. (2024). Additionally, a simple yet effective multimodal and multilingual method for parsing semi-structured forms is presented in Cheng et al. (2025). Despite the variety of models proposed, comparative studies evaluating text-only, multimodal, and layout-aware LLMs remain limited Van Landeghem et al. (2023); Stanisławek et al. (2021).
Preprocessing, Chunking, Prompting, Postprocessing, and Evaluation Techniques. Chain-of-Thought (CoT) prompting Wei et al. (2022) enhances reasoning in complex LRD extraction tasks, while diverse prompt-response datasets Zmigrod et al. (2024) improve LLM robustness. Instruction-finetuned LLMs have also demonstrated effectiveness in domain-specific applications, such as clinical and biomedical tasks, where zero-shot and few-shot enable adaptive extraction without extensive fine-tuning Labrak et al. (2024). Postprocessing techniques, including text normalization and schema alignment Fürst et al. (2023).
@@ Acknowledgments
Our experimental evaluation focused on the Visually Rich Document Understanding (VRDU) dataset Wang et al. (2023b) and the FUNSD dataset Jaume et al. (2019). VRDU was chosen for its comprehensive OCR outputs, well-defined schema, and diverse template types (Amendment, Short-Form, and Dissemination) that capture complex layouts. It also supports generalization tasks–Single Template Learning (STL), Unseen Template Learning (UTL), and Mixed Template Learning (MTL)–making it suitable for visually rich document understanding. In contrast, FUNSD was selected due to its widespread use in document understanding tasks, particularly for extracting structured information from simpler, form-based layouts. While FUNSD does not share the layout diversity and complexity of VRDU, it serves as a useful baseline for form-centric document analysis. This provides a different perspective by offering insights into scenarios where key-value pairs are organized within relatively simple templates.